OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (–1, 5) and has a slope of –4? y = x +

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Demonx341

TheSmartOne (thesmartone):

y=mx+b y=-4x+b Plug in (-1,5) and solve for b And then just plug back in b

OpenStudy (anonymous):

@lsherron this is my last question -_-

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

Well. Slope intercept is y= mx + b. With B being the y intercept. If it contains 1,1 and the slope is positive 6, do the equation 1 = 6 (1) + b. 1 = 6 + b -5 = b. So the answer is y = 6x - 5

OpenStudy (studygurl14):

Do what @TheSmartOne said @RainbowBubbles01

TheSmartOne (thesmartone):

Because @lsherron 's answer is wrong...

OpenStudy (studygurl14):

@Isherron "has a slope of –4", not 6

OpenStudy (studygurl14):

"has a slope of -4"

OpenStudy (studygurl14):

So... @TheSmartOne is correct

OpenStudy (anonymous):

so what was it?

OpenStudy (anonymous):

nvm i passed

OpenStudy (anonymous):

and it was my samester test prt.1

TheSmartOne (thesmartone):

y=-4x+1 was it.. Good job on passing.